What volume of 12.0 M sulfuric acid would be required to prepare 250 mL of 6.5 M sulfuric acid?
1
Expert's answer
2019-09-17T04:58:00-0400

c1(H2SO4) = 12.0M

c2(H2SO4) = 6.5M

V2(H2SO4) = 250 mL = 0.25 L

V1(H2SO4) - ?


The amount of sulfuric acid (in moles) in the required volume of stock solution is the same as in the diluted sample:

n1(H2SO4) = n2(H2SO4)

c1(H2SO4) V1(H2SO4) = c2(H2SO4) V2(H2SO4)


"V_1(H_2SO_4)=\frac{c_2(H_2SO_4) \cdot V_2(H_2SO_4)}{c_1(H_2SO_4)}=\frac{6.5 \frac{mol}{L} \cdot 0.25 L}{12.0 \frac{mol}{L}}=0.135 L = 135 mL"